Rwkit:

Consider a car with a cash price of $4,000, offered on the installment plan for $146 per month for 36 months. Calculate (a) the total installment price, (b) the carrying charges, and (c) the number of months needed to save the money at the monthly rate to buy the car for its cash price. If applicable use a thousands comma. a. $ b. $ c. months

Serenity1jackson:

um it is not A

XxAngelGirlxX:

carrying charges = 146 * 36 - 4000 = ? 146* 36= then subtract that answer from 4000 to find your answer

Rwkit:

Its three parts to this question a b and c

XxAngelGirlxX:

yes 4000 would be the first

XxAngelGirlxX:

then whatever your answer is from my first question is your second

Serenity1jackson:

yeah ^^^^ he are she is right

XxAngelGirlxX:

she lol

XxAngelGirlxX:

so whats the second part answer?

Serenity1jackson:

um you asking me

XxAngelGirlxX:

no, im asking @Rwkit

Rwkit:

Let me think

XxAngelGirlxX:

oki

Rwkit:

I think its 4$ not sure

XxAngelGirlxX:

A. 4,000 b. (146) rounding up, 28 months

Rwkit:

Thanks